OpenStudy (anonymous):

Solve by the elimination method: x + y = 18 8x - y = 63

OpenStudy (vishweshshrimali5):

I can help you

OpenStudy (anonymous):

Thanks. thats nice of you

OpenStudy (vishweshshrimali5):

See, Just add both the equations

OpenStudy (vishweshshrimali5):

then you will get 9x = 81 => x = 9

OpenStudy (vishweshshrimali5):

This is what you needed !!

OpenStudy (vishweshshrimali5):

Are you satisfied with the answer ?

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

y=18-x 8x-(18-x)=63 8x-18+x=63 9x=63+18 9x=81 x=9 if i made no mistake

OpenStudy (vishweshshrimali5):

Yaa you are right too racht

OpenStudy (anonymous):

eliminate one variable from the first equation, them apply it into the second. then solve the new equation by moving all the unknown to the same side, the known to the other side

OpenStudy (anonymous):

You got it right. thanks so much

OpenStudy (vishweshshrimali5):

Hey congratss RachT you got a medal !!!! :)

OpenStudy (vishweshshrimali5):

Hey I got a medal tooo hurray !

OpenStudy (vishweshshrimali5):

Hey kuoldan you got a medal too